OpenStudy (anonymous):

If tanya rides her bike at speed 15mi/h, how far will she ride in 2 hours?

OpenStudy (kyosuke):

V = d * t then 15 = d * 2 distance = 7.5 Millas Luck.! Kyosuke =)

OpenStudy (whpalmer4):

\[d = rt\] where \(d\) is distance traveled, \(r\) is rate (or speed) and \(t\) is time. Tanya's rate is 15 mi/h, and the time is 2 h. Therefore\[d=(15 mi/h)(2h) = 30 mi. Note that @kyosuke could have spotted the error if the units had been included in the calculation instead of added on at the end: miles * hours is not miles/hour. If the units don't have the right form, the answer will be incorrect.

OpenStudy (kyosuke):

sorry please ,I'm wrong. correct D = v * t entonces d = 15* 2 = distancia 30 Millas.! Kyosuke =)